3x-5(3x-7)=2(x+9)+45

Answer by stanbon(75887)

3x-5(3x-7)=2(x+9)+45
3x-15x+35 = 2x+18+45
-12x + 35 = 2x+ 63
14x = - 28
x = -2
============
Cheers,
Stan H.